You Have to answer 3 essay questions for an exam . There are 5 essays to choose from . How many different groups of 3 essays could you possibly choose ?
Answer by jim_thompson5910(35256)

If order mattered, then you'd have 5*4*3 = 20*3 = 60 different possible permutations

However, order doesn't matter. So there are 60/(3!) = 60/(3*2*1) = 60/6 = 10 different combinations possible.

The final answer is 10.